OpenStudy (anonymous):

change this equation into slope-intercept form –x + 3y = 9

OpenStudy (lgbasallote):

hint (i'll give you a demonstration) if you have something like 2x + 3y = 8 simply isolate y \[\implies 3y = -2y + 8\] \[\implies y = \frac{-2y + 8}{3}\] does that help?

OpenStudy (anonymous):

what happens to the x though?

OpenStudy (lgbasallote):

uhhh that should be x not y..*facepalm* \[\implies 3y = -2x + 8\] \[\implies y = \frac{-2x + 8}{3}\] sorry bout that

OpenStudy (anonymous):

sorry i was confused . lol thankyou :))